### v1

@@ Abstract
Accurate and reliable forecasting models are critical for guiding public health responses and policy decisions during pandemics such as COVID-19. Retrospective evaluation of model performance is essential for improving epidemic forecasting capabilities. In this study, we used COVID-19 wastewater data from CDC’s National Wastewater Surveillance System to generate sequential weekly retrospective forecasts for the United States from March 2022 through September 2024, both at the national level and for four major regions (Northeast, Midwest, South, and West). We produced 133 weekly forecasts using 11 models, including ARIMA, generalized additive models (GAM), simple linear regression (SLR), Prophet, and the $n$-sub-epidemic framework (top-ranked, weighted-ensemble, and unweighted-ensemble variants). Forecast performance was assessed using mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% prediction interval coverage. The $n$-sub-epidemic unweighted ensembles outperformed all other models at 3–4-week horizons, particularly at the national level and in the Midwest and West. ARIMA and GAM performed best at 1–2-week horizons in most regions, whereas Prophet and SLR consistently underperformed across regions and horizons. These findings highlight the value of region-specific modeling strategies and demonstrate the utility of the $n$-sub-epidemic framework for real-time outbreak forecasting using wastewater surveillance data.
@@ 1 Introduction
The COVID-19 pandemic has highlighted the urgent need for efficient, scalable, and timely disease surveillance methods. Traditional epidemiological approaches rely primarily on clinical testing and hospital records, which, while valuable, are often constrained by delays in reporting, limited testing availability, and the under-detection of asymptomatic infections (nixon2022real; alvarez2023limitations). These limitations have underscored the importance of alternative surveillance methods, especially wastewater-based epidemiology (WBE), which has emerged as a cost-effective, non-invasive, and real-time tool to monitor SARS-CoV-2 transmission at the community level. Viral particles are shed in human waste even before individuals exhibit symptoms or seek medical attention (medema2020presence; kitajima2020sars). Unlike traditional testing, which depends on individual participation and healthcare access, WBE provides a comprehensive snapshot of community-wide infection levels. As a result, it has been widely implemented across multiple countries to track and forecast COVID-19 trends (peccia2020measurement). Its potential for early detection makes WBE especially valuable for improving calibration and lead-time accuracy in epidemic forecasting.
@@
Recent research has demonstrated a strong and consistent correlation between SARS-CoV-2 RNA concentrations in wastewater and the number of reported COVID-19 cases (li2023correlation). This relationship establishes WBE as an effective early warning system, allowing public health officials to detect emerging outbreaks and monitor infection trends  (gonzalez2020covid). WBE has been successfully applied to epidemic forecasting on different geographical scales, from localized wastewater treatment plants to national surveillance networks (joseph2022assessing; daza2023model). For instance, a study in Catalonia (Spain) demonstrated the potential of WBE to predict SARS-CoV-2 incidence across multiple regions, reinforcing its role in epidemiological forecasting (joseph2022assessing). Similarly, research in the United States (U.S.) showed that wastewater data could predict weekly COVID-19 hospital admissions up to four weeks in advance, emphasizing its value in public health preparedness (li2023wastewater). In rural communities where clinical testing is often limited, WBE has proven particularly beneficial. In a recent multi-state analysis of 189 wastewater treatment plants during 2023-2024, it was demonstrated that SARS-CoV-2 RNA concentrations precede hospitalization admissions by 2 to 12 days, with a lead time varying by state-level WBE coverage and facility characteristics (schenk2024sars). A study in Idaho successfully integrated wastewater data with SEIR modeling to forecast outbreaks with lead times of up to 11 days, showcasing its ability to enhance surveillance in underserved areas (meadows2025epidemiological). However, most of these studies have focused on short-term forecasts or single geographic areas, leaving a gap in understanding long-term WBE forecasting performance across diverse U.S. regions. With the gradual de-escalation of widespread clinical testing and case-based reporting in the U.S., wastewater surveillance is becoming an increasingly critical data source for public health forecasting. This shift further underscores the need to evaluate the forecasting performance of models applied to wastewater data across different geographic and temporal contexts.
@@
In this study, we conducted a comprehensive retrospective evaluation of 1- to 4-week forecast performance across 11 different models. Our analysis includes n-sub-epidemic models alongside statistical approaches such as autoregressive integrated moving average (ARIMA), generalized additive models (GAMs), Simple Linear Regression (SLR), and Meta (Facebook’s) Prophet model. We assess forecasting performance at both national and regional levels, covering four U.S. regions: the Midwest, Northeast, South, and West. The evaluation period spans from March 5, 2022, to September 14, 2024. To ensure a rigorous and standardized assessment, we utilize key epidemic forecasting performance metrics, including mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS). Integrating dynamic modeling approaches with wastewater surveillance data has been proposed as a promising strategy to enhance predictive accuracy and real-time outbreak monitoring. Our findings provide insights into the optimal selection of forecasting models for wastewater-based epidemic surveillance across different geographic contexts and time horizons.
@@ 2.2 Model calibration and specification
We used a total of 11 models in our analysis, including the top-ranked models, ensemble models (weighted and unweighted) from the n-sub-epidemic framework, along with ARIMA, GAM, GLM, and SLR (each estimated under a normality assumption), and the Prophet model. A detailed description of the methodologies for these models is provided below.
The CDC publishes wastewater data weekly. For each model and location in our analysis, we used all data between January 1, 2022, and September 14, 2024, released by the CDC on October 26, 2024, to construct rolling calibration windows. The forecasting period began on March 5, 2022 (10 weeks after the first considered data point), and advanced in one-week increments, with models refit each week using the most recent 10 weeks of data to generate 1- to 4-week-ahead forecasts; the 10-week calibration window balances sufficient data for capturing recent dynamics while minimizing the influence of changing variability and structural breaks, consistent with short-term epidemic forecasting practice (bleichrodt2024evaluating; bleichrodt2023real).
We assumed a normal error structure in the observed data. This decision was based on two main considerations: (i) the short 10-week calibration window helped minimize the influence of changing variability in the data over time, and (ii) the models demonstrated consistently high calibration coverage, suggesting that the normality assumption provided an adequate approximation of the residual distribution over the fitting period.
@@ 4 Results
Our retrospective forecasting study generated 133 weekly forecast sets spanning March 2022 through September 2024, producing over 10,000 individual forecast-observation pairs across all regions, models, and horizons. Figures 4 through 8 present aggregated performance metrics (mean values computed across all 133 forecast weeks) to identify systematic patterns in model performance. To complement these averages, we also present the full distributions of MAE, MSE, and WIS across forecast weeks in Supplementary Figures S1–S5 (boxplots for each region), which highlight variability and consistency of model performance over time. This aggregation approach balances comprehensive evaluation with interpretability. Throughout the Results, we reference models using the nomenclature established in Section 2.8: Rank 1-3 (top n-sub-epidemic models by AICc), EM2 W/UW and EM3 W/UW (2- and 3-model ensembles, weighted and unweighted), and comparison models ARIMA, GAM, SLR, and Prophet.
@@ 4.1 National
To help evaluate the magnitude and variability of these differences, we present the distributions of all metrics across models, forecast periods, and horizons at the national level in Supplementary Fig. S1 (with corresponding regional distributions in Supplementary Figs. S2–S5). Performance metric distributions across all 133 forecast periods (Supplementary Figs. S1–S5) reveal the consistency of model rankings. Focusing on the national level, Supplementary Fig. S1 shows that GAM and EM3 UW consistently achieve the lowest MAE values among all models, with their distributions clearly separated from SLR and Prophet. The boxplots show partial overlap among the top-performing models (GAM, EM3 UW, ARIMA, Rank 3), indicating that while ensemble approaches offer reliable advantages, performance differences among leading models vary across forecast periods. This pattern suggests that the ensemble approach provides dependable, though not universal, improvements across changing epidemic dynamics.
Skill scores relative to the SLR baseline (Section 3.5) quantify forecast improvement beyond simple trend extrapolation. At the national level, EM3 UW achieves MAE skill scores of 46% at 1-week ahead and 40–41% at 2–4 week horizons, representing substantial error reduction compared to linear trend forecasting (Supplementary Figure S6). For MSE, skill scores reach 55–70%, while WIS skill scores range from 49–54% across all horizons. These consistently strong values demonstrate that ensemble forecasting provides meaningful improvements over naive baseline methods throughout the forecast window. Regional skill score patterns (Supplementary Figures S7–S10) show similar results, with ensemble models consistently achieving positive skill scores across all regions and horizons.
@@ 4.2 Midwest
Results for the Midwest region are summarized in Fig. 5. Forecast errors increase with horizon length, but overall model rankings remain relatively stable. At the 1-week horizon, GAM and ARIMA are among the best-performing models, with GAM attaining the lowest MAE (0.63) and the lowest MSE (0.96), while ARIMA also performs competitively, particularly for MAE (0.64) and WIS (0.47). The EM3 UW ensemble provides the highest 95% PI coverage (87.97%), indicating strong uncertainty quantification even for near-term forecasts.
@@
Figure 5: Log-transformed averages of MAE, MSE, and WIS, and the (untransformed) average 95% PI coverage across all models for 1–4 week horizons in the Midwest region, spanning the period from March 5, 2022, to September 14, 2024 (133 forecasts). Lighter shades indicate smaller (better) values for MAE, MSE, and WIS, whereas darker shades indicate larger errors. For coverage, values closer to 95% indicate better performance. Black cells highlight the best-performing model(s) for each forecast horizon and metric.
@@ 4.3 Northeast
Fig. 6 summarizes the results for the Northeast region. Across all horizons, this region shows some of the highest MAE, MSE, and WIS values, indicating that the Northeast was among the most challenging settings for forecasting in our study. We discuss potential reasons for this in detail in Section 5.1.
@@
Figure 6: Log-transformed averages of MAE, MSE, and WIS, and the (untransformed) average 95% PI coverage across all models for 1–4 week horizons in the Northeast region, spanning the period from March 5, 2022, to September 14, 2024 (133 forecasts). Lighter shades indicate smaller (better) values for MAE, MSE, and WIS, whereas darker shades indicate larger errors. For coverage, values closer to 95% indicate better performance. Black cells highlight the best-performing model(s) for each forecast horizon and metric.
@@ 4.4 South
Fig. 7 summarizes the results for the South region. This region exhibits moderate error levels relative to the Northeast and Midwest, with clearer separation among models and more stable performance patterns across horizons. Overall, ARIMA and EM3 UW consistently rank among the best-performing models across all forecast horizons, with GAM also performing well at 1–2 weeks.
@@
Figure 7: Log-transformed averages of MAE, MSE, and WIS, and the (untransformed) average 95% PI coverage across all models for 1–4 week horizons in the South region, spanning the period from March 5, 2022, to September 14, 2024 (133 forecasts). Lighter shades indicate smaller (better) values for MAE, MSE, and WIS, whereas darker shades indicate larger errors. For coverage, values closer to 95% indicate better performance. Black cells highlight the best-performing model(s) for each forecast horizon and metric.
@@ 4.5 West
Fig. 8 presents the results for the West region. There is clear separation among models and strong performance from both ARIMA and the unweighted n-sub-epidemic ensembles, particularly EM3–UW. Overall, these models consistently rank among the top-performing approaches across all forecast horizons.
@@
Figure 8: Log-transformed averages of MAE, MSE, and WIS, and the (untransformed) average 95% PI coverage across all models for 1–4 week horizons in the West region, spanning the period from March 5, 2022, to September 14, 2024 (133 forecasts). Lighter shades indicate smaller (better) values for MAE, MSE, and WIS, whereas darker shades indicate larger errors. For coverage, values closer to 95% indicate better performance. Black cells highlight the best-performing model(s) for each forecast horizon and metric.
@@ 5 Discussion
Model performance varied across regions and forecast horizons. For example, EM3 UW dominated in the Midwest, especially at 3- to 4-week forecast horizons, while GAM and ARIMA excelled at 1- to 2-week-ahead forecasts. In the South and Northeast regions, ARIMA consistently performed well in the short term (1- to 2-week), but ensemble models like EM3 UW and EM2 UW showed more substantial 3- to 4-week reliability. In the West, EM2 UW and EM3 UW performed very well at 3- to 4-week forecasts. These results emphasize the need for context-specific modeling strategies, particularly when geographic, demographic, and reporting differences can impact model performance (pei2020differential).

### v2

@@ Abstract
Accurate and reliable forecasting models are critical for guiding public health responses and policy decisions during pandemics such as COVID-19. Retrospective evaluation of forecasting performance provides an essential framework for assessing and improving epidemic prediction methods. In this study, we used COVID-19 wastewater data from CDC’s National Wastewater Surveillance System to generate sequential weekly retrospective out-of-sample forecasts for the United States from March 2022 through September 2024, both at the national level and for four major regions (Northeast, Midwest, South, and West). We produced 133 weekly forecasts using 11 models, including ARIMA, generalized additive models (GAM), simple linear regression (SLR), Prophet, and the $n$-sub-epidemic framework (top-ranked, weighted-ensemble, and unweighted-ensemble variants). Forecast performance was assessed using mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% prediction interval coverage. The $n$-sub-epidemic unweighted ensembles outperformed all other models at 3–4-week horizons, particularly at the national level and in the Midwest and West. ARIMA and GAM performed best at 1–2-week horizons in most regions, whereas Prophet and SLR consistently underperformed across regions and horizons. These findings highlight the value of region-specific modeling strategies and demonstrate the utility of the $n$-sub-epidemic framework for real-time outbreak forecasting using wastewater surveillance data.
@@ 1 Introduction
The COVID-19 pandemic has highlighted the urgent need for efficient, scalable, and timely disease surveillance methods. Traditional epidemiological approaches rely primarily on clinical testing and hospital records, which, while valuable, are often constrained by delays in reporting, limited testing availability, and the under-detection of asymptomatic infections (nixon2022real; alvarez2023limitations). These limitations have underscored the importance of alternative surveillance methods, especially wastewater-based epidemiology (WBE), which has emerged as a cost-effective, non-invasive, and real-time tool to monitor SARS-CoV-2 transmission at the community level (medema2020presence; kitajima2020sars; parkins2024wastewater; rashid2024scoping). Viral particles are shed in human waste even before individuals exhibit symptoms or seek medical attention (medema2020presence; kitajima2020sars). Unlike traditional testing, which depends on individual participation and healthcare access, WBE provides a comprehensive snapshot of community-wide infection levels. As a result, it has been implemented in a wide range of settings internationally to track SARS-CoV-2 dynamics (rashid2024scoping; parkins2024wastewater). Because wastewater signals often precede clinical indicators, WBE has increasingly been explored as a complementary data stream to support epidemic monitoring and short-term forecasting efforts.
@@
Recent research has demonstrated that SARS-CoV-2 RNA concentrations in wastewater are often associated with reported COVID-19 cases, hospitalizations, and other epidemiologic indicators (peccia2020measurement; weidhaas2021correlation; li2023correlation). This relationship establishes WBE as an effective early warning system, allowing public health officials to detect emerging outbreaks and monitor infection trends (gonzalez2020covid). At the same time, the strength and timing of these relationships are not uniform across locations or time periods. Systematic review evidence has shown substantial variability in wastewater–clinical correlations across studies, with estimates influenced by catchment characteristics, sampling frequency, environmental conditions, and clinical testing coverage (li2023correlation). In addition, these relationships may vary across dominant variant periods and changes in healthcare-seeking or testing behavior (zhan2023correlative). WBE has been successfully applied to epidemic forecasting on different geographical scales, from localized wastewater treatment plants to national surveillance networks (joseph2022assessing; li2023wastewater). For instance, a study in Catalonia (Spain) demonstrated the potential of WBE to predict SARS-CoV-2 incidence across multiple regions, reinforcing its role in epidemiological forecasting (joseph2022assessing). Similarly, research in the United States (U.S.) showed that wastewater data could predict weekly COVID-19 hospital admissions up to four weeks in advance, emphasizing its value in public health preparedness (li2023wastewater). In rural communities where clinical testing is often limited, WBE has proven particularly beneficial. In a recent multi-state analysis of 189 wastewater treatment plants during 2023–2024, it was demonstrated that SARS-CoV-2 RNA concentrations precede hospitalization admissions by 2 to 12 days, with lead time varying by state-level WBE coverage and facility characteristics (schenk2024sars). A study in Idaho successfully integrated wastewater data with susceptible–exposed–infectious–recovered (SEIR) modeling to forecast outbreaks with lead times of up to 11 days, showcasing its ability to enhance surveillance in underserved areas (meadows2025epidemiological). However, most of these studies have focused on short-term forecasts or single geographic areas, leaving a gap in understanding the comparative performance of forecasting models applied retrospectively across diverse U.S. regions. With the gradual de-escalation of widespread clinical testing and case-based reporting in the U.S., wastewater epidemiology is becoming an increasingly important data source for forecasting potential public health impacts (nixon2022real; parkins2024wastewater). This shift further underscores the need to evaluate the forecasting performance of models applied to wastewater data across different geographic and temporal contexts.
@@
In this study, we conducted a comprehensive retrospective out-of-sample evaluation of 1- to 4-week forecast performance across 11 different models. Our analysis includes n-sub-epidemic models alongside statistical approaches such as autoregressive integrated moving average (ARIMA), generalized additive models (GAMs), generalized logistic growth models (GLMs), simple linear regression (SLR), and Meta (Facebook’s) Prophet model. We assess forecasting performance at both national and regional levels, covering four U.S. regions: the Midwest, Northeast, South, and West. The evaluation period spans from March 5, 2022, to September 14, 2024. To ensure a rigorous and standardized assessment, we utilize key epidemic forecasting performance metrics, including mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS). By systematically comparing multiple forecasting approaches using a rolling-origin retrospective framework, our study provides insights into the relative strengths of statistical and phenomenological models for short-term epidemic forecasting using wastewater surveillance data.
@@ 2.2 Model calibration and specification
We used a total of 11 models in our analysis, including the top-ranked models and ensemble models (weighted and unweighted) from the n-sub-epidemic framework, together with autoregressive integrated moving average (ARIMA), generalized additive models (GAMs), generalized logistic growth models (GLMs), simple linear regression (SLR), and Prophet models. A detailed description of these models is provided below.
We analyzed the CDC National Wastewater Surveillance System (NWSS) weekly wastewater viral activity level (WVAL) time series available in the October 26, 2024 CDC release at the national level and for four U.S. regions. In this study, “all data” refers to the complete sequence of weekly published WVAL values from January 1, 2022, through September 14, 2024, which was used only to define the retrospective sequence of forecast origins and evaluation targets. We used the published weekly national and regional WVAL series as provided in that CDC/NWSS release and did not reconstruct these series from raw site-level wastewater measurements. No additional rolling average, interpolation, or site-level reweighting was applied.
For each forecast origin, models were fit only to the most recent 10 weeks of data available at that time. Starting on March 5, 2022 (10 weeks after the first observation in the study period), we advanced the forecast origin by one week at a time, re-estimated all models using the rolling calibration window [t9,t][t-9,t], and generated 1- to 4-week-ahead forecasts for weeks $t+1$ through $t+4$. Forecast accuracy was then evaluated against the subsequently observed WVAL values. Thus, the full January 2022–September 2024 record was used only to define the sequence of retrospective forecast origins and out-of-sample evaluation targets; no future observations were used in model fitting at any forecast date, and training and evaluation data were kept separate throughout.
Because our analysis was conducted on the published weekly WVAL series rather than raw site-level wastewater measurements, differences in sampling frequency across contributing sites or jurisdictions were not modeled explicitly as separate weights in our forecasting framework. Accordingly, our comparison evaluates model performance on the published weekly surveillance signal rather than on a site-level harmonized dataset. This is an important scope condition and should be considered when interpreting differences in forecast performance across regions and time periods.
The 10-week calibration window was selected to balance having sufficient observations to capture recent local dynamics while limiting the influence of older observations, changing variability, and possible structural breaks in the wastewater signal, consistent with prior short-term epidemic forecasting studies (bleichrodt2024evaluating; bleichrodt2023real).
All comparison models used only historical WVAL values within the calibration window and did not incorporate external covariates. The n-sub-epidemic framework assumes that the observed trajectory can be represented as a superposition of overlapping sub-epidemic waves. ARIMA assumes that short-term dependence in the differenced series can be captured through autoregressive and moving-average components. GAM assume a smooth nonlinear relationship between WVAL and time over the calibration window. GLM assume generalized logistic growth dynamics over the fitting period. SLR assumes an approximately linear trend in WVAL over the calibration window. Prophet assumes that the time series can be represented through decomposable trend components estimated from recent observations.
We assumed a normal error structure in the observed data for the models estimated under this framework. This decision was based on two considerations: (i) the short 10-week calibration window helped reduce the influence of time-varying variance and structural changes, and (ii) the models demonstrated high empirical calibration coverage, suggesting that the normal approximation was adequate over the fitting period.
@@ 4 Results
Using a rolling-origin retrospective evaluation design, we generated 133 out-of-sample weekly forecast sets spanning March 5, 2022, through September 14, 2024, producing over 10,000 forecast-observation pairs across all regions, models, and forecast horizons. At each forecast origin, models were fit only to the preceding 10 weeks of WVAL data and then evaluated against subsequently observed values at 1- to 4-week horizons. Figures 4, Supplementary Figures S2, S4, S6, and S8 present aggregated performance metrics (mean values computed across all 133 forecast weeks) to identify systematic patterns in model performance. To complement these averages, we also present the full distributions of MAE, MSE, and WIS across forecast weeks in Supplementary Figures S1–S5 (boxplots for each region), which highlight variability and consistency of model performance over time. This aggregation approach balances comprehensive evaluation with interpretability. Throughout the Results, we reference models using the nomenclature established in Section 2.8: Rank 1-3 (top n-sub-epidemic models by AICc), EM2 W/UW and EM3 W/UW (2- and 3-model ensembles, weighted and unweighted), and comparison models ARIMA, GAM, SLR, and Prophet.
@@ 4.1 National
To help evaluate the magnitude and variability of these differences, we present the distributions of all metrics across models, forecast periods, and horizons at the national level in Supplementary Fig. S1 (with corresponding regional distributions in Supplementary Figs. S3, S5, S7, and S9). Performance metric distributions across all 133 forecast periods reveal consistency of model rankings. Focusing on the national level, Supplementary Fig. S1 shows that GAM and EM3 UW consistently achieve the lowest MAE values among all models, with their distributions clearly separated from SLR and Prophet. The boxplots show partial overlap among the top-performing models (GAM, EM3 UW, ARIMA, Rank 3), indicating that while ensemble approaches offer reliable advantages, performance differences among leading models vary across forecast periods. This pattern suggests that the ensemble approach provides dependable, though not universal, improvements across changing epidemic dynamics.
Skill scores relative to the SLR baseline (Section 3.5) quantify forecast improvement beyond simple trend extrapolation. At the national level, EM3 UW achieves MAE skill scores of 46% at 1-week ahead and 40–41% at 2–4 week horizons, representing substantial error reduction compared to linear trend forecasting (Supplementary Figure S 10). For MSE, skill scores reach 55–70%, while WIS skill scores range from 49–54% across all horizons. These consistently strong values demonstrate that ensemble forecasting provides meaningful improvements over naive baseline methods throughout the forecast window. Regional skill score patterns (Supplementary Figures S11–S14) show similar results, with ensemble models consistently achieving positive skill scores across all regions and horizons.
@@ 4.2 Midwest
Results for the Midwest region are summarized in supplementary Fig. 2. Forecast errors increase with horizon length, but overall model rankings remain relatively stable. At the 1-week horizon, GAM and ARIMA are among the best-performing models, with GAM attaining the lowest MAE (0.63) and the lowest MSE (0.96), while ARIMA also performs competitively, particularly for MAE (0.64) and WIS (0.47). The EM3 UW ensemble provides the highest 95% PI coverage (87.97%), indicating strong uncertainty quantification even for near-term forecasts.
@@ 4.3 Northeast
Supplementary Fig. S4 summarizes the results for the Northeast region. Across all horizons, this region shows some of the highest MAE, MSE, and WIS values, indicating that the Northeast was among the most challenging settings for forecasting in our study. We discuss potential reasons for this in detail in Section 5.1.
@@ 4.4 South
Supplementary Fig. S6 summarizes the results for the South region. This region exhibits moderate error levels relative to the Northeast and Midwest, with clearer separation among models and more stable performance patterns across horizons. Overall, ARIMA and EM3 UW consistently rank among the best-performing models across all forecast horizons, with GAM also performing well at 1–2 weeks.
@@ 4.5 West
Supplementary Fig. S8 presents the results for the West region. There is clear separation among models and strong performance from both ARIMA and the unweighted n-sub-epidemic ensembles, particularly EM3–UW. Overall, these models consistently rank among the top-performing approaches across all forecast horizons.
@@ 5 Discussion
Model performance varied across regions and forecast horizons. For example, EM3 UW dominated in the Midwest, especially at 3- to 4-week forecast horizons, while GAM and ARIMA excelled at 1- to 2-week-ahead forecasts. In the South and Northeast regions, ARIMA consistently performed well in the short term (1- to 2-week), but ensemble models like EM3 UW and EM2 UW showed more substantial 3- to 4-week reliability. In the West, EM2 UW and EM3 UW performed very well at 3- to 4-week forecasts. These results emphasize the need for context-specific modeling strategies, but the observed heterogeneity should be interpreted cautiously because regional differences in forecast performance may reflect both model behavior and differences in the wastewater surveillance signal itself. Prior wastewater studies have shown that signal interpretation can vary with catchment characteristics, sampling strategy, environmental conditions, and the relationship between wastewater measurements and epidemiologic indicators. (bertels2022factors; zhan2023correlative)